OpenStudy (zacharydoc):

Solve for x: 3x − 5 = 2x + 6. 1 −1 11 −11

OpenStudy (kenrick143):

Transposing getting 3x-2x=5+6 x=11

OpenStudy (zacharydoc):

thankyou

OpenStudy (welshfella):

the aim is to get all terms in x on ine side of the equation and all the numbers on the other side.

OpenStudy (welshfella):

pleas dont give direct answers . Try and involve the questioner in the process of explaining.

OpenStudy (kenrick143):

Sorry..